OpenStudy (anonymous):

Divide 36x3 – 96x2 + 48x by 12x. Answer: A) 3x2 – 8x + 4 B) 3x3 + 8x – 4 C) 3x2 + 8x – 4 D) 3x3 – 8x + 4 help please:)

OpenStudy (anonymous):

All terms contain the common factor 12x. So how many times does it divide into each?\[\Large \frac{{36{x^3}}}{{12x}} = 3{x^2}\]CAn you work out the rest?

OpenStudy (anonymous):

36x3/12x=3x2 -96x2/12x=-8x 48x/12x=4 so 3x2 -8x + 4

OpenStudy (anonymous):

yes, thanks guys:))

OpenStudy (anonymous):

Otherwise 12x(3x^2-8x=4)/12x

OpenStudy (anonymous):

sorry 12x(3x^2-8x+4)/12x